If Xo = 3 and Xn = (5xn-1 + 7) mod 200 find x], ..., X 10.
Answer #1

zos (520-1 +7 ) mod 200. |_ Given 20 = 3. 5 x, = (5x+7) mod 200 x1 = (5.3+7) mod 200. æ, - 22cmod 200) #2 = (5%,+7) mod 200 -23. = 592 (mod 200) - [ 2x200 + 192) mod 200. x3 192 (mod 200) 24 = (5x3 + 7) (mod 200) = (967) mod 200. = (167) mod 200 X5 E